The value of a piece of factory equipment after 3 years of use is V=100(0.5)x where X is a random variable having moment generating function MX(t)=(1-2t)-1,t<1/2. Calculate the expected value of this piece of equipment after three years of use.


1
Expert's answer
2021-07-22T06:09:37-0400

By the problem the value of a piece of factory equipment after three years of use is "V=100(0.5)^X" where X is a random variable having moment generating function

"M_X(t)=E(e^{tX})=\frac{1}{(1-2t)}, for t<\frac{1}{2}."

So the expected value of this piece of equipment after three years of use

"E(V) = E(100(0.5)^X) = 100E(0.5^X) \\

E(V) = 100E(e^{ln0.5^X})=100E(e^{Xln0.5})=100M_X(ln0.5) \\

ln(0.5)= -0.6931 <0.5 \\

E(V) = \frac{100}{1-2 \times (-0.6931)}= 41.91"